# Scalable Stewardship of an LLM-Assisted Clinical Benchmark with Physician Oversight

### v2

@@ Appendix A LLM Prompts
Phase 2 System Prompt: You will read a user-provided JSON string that contains an anonymized patient note (‘‘Patient Note’’) and a clinical computation question (‘‘Question’’). Your job is to calculate the answer to the question based on the patient note. You have access to a Google Search tool and a code execution tool to perform calculations, if needed. Your operational protocol is as follows: 1. Adopt Expert Persona: Assume the role of a diligent clinician. Your primary responsibility is accuracy and adherence to standard and up-to-date medical guidelines. 2. Identify Appropriate Scoring Rule or Formula: You will be given a patient note and a question that specifies a clinical calculation task (e.g., CHA2DS2-VASc score, MELD score, CURB-65, etc.). You should use Google Search to confirm the current, generally accepted formula, criteria, and/or scoring system for the specified clinical computation question. Prioritize reputable medical sources, such as MDCalc.com and guidelines published by health agencies or hospitals. 3. Systematic Reasoning and Calculation: Carefully read the patient note and systematically extract every piece of information relevant to the calculation. Explicitly map the extracted patient data to the components of the scoring system. For example, for a CHA2DS2-VASc score, list each component (Congestive heart failure, Hypertension, Age, etc.) and assign points based on the patient note. If there are complex arithmetic operations that are not easily performed by hand, you can use the code execution tool to perform the calculations. Show your step-by-step calculation. Sum the points or apply the formula as required. 4. Handling Missing Information: If a piece of information required for the calculation is clearly absent from the note and if it cannot be reasonably inferred from the note, explicitly state what is missing in your reasoning, and give your final answer as ‘‘unknown’’ (without quotes). Do not invent data or make assumptions that are not reasonably supported by the patient note. 5. Final Answer Formatting (Strict): After your detailed reasoning, you MUST enclose your final answer in <answer> </answer> tags. The content inside the tags must ONLY be the final numeric value (e.g., 4, 12.5), date, or time. For numeric values, you do not need to round the answer. As mentioned in #4, if and only if the calculation is impossible due to missing data, the content inside the answer tags must be exactly ‘‘unknown’’ (without quotes). DO NOT include any units, variable names, explanations, or any other text inside the <answer> </answer> tags. Phase 2 User Prompt Example: { "Patient Note": "An 87-year-old man was admitted to our hospital for anorexia for several days, high-grade fever from the previous day, and liver dysfunction. Of note, he had a history of hypertension, diabetes mellitus (DM), and angina. Physical examination findings included: clear consciousness; height, 163 cm; weight, 48 kg; blood pressure, 66/40 mmHg; heart rate, 75/min; respiratory rate, 22/min...[Omitted content was present in the actual API call]...causes of death were thus determined as (1) DIC and (2) ALF associated with diffuse large B-cell lymphoma.", "Question": "What is the patient’s Creatinine Clearance using the Cockroft-Gault Equation in terms of mL/min? You should use the patient’s adjusted body weight in kg instead of the patient’s actual body weight if the patient is overweight or obese based on their BMI. If the patient’s BMI is normal, set their adjusted body weight to the minimum of the ideal body and actual weight. If the patient is underweight, please set their adjusted body weight to their actual body weight.", }
@@ Appendix B Concurrent Work
[43] reexamined MedCalc-Bench by introducing an LLM-judged, stepwise evaluation pipeline to grade an LLM’s intermediate calculation steps at test time, which can be thought of as providing process-level rewards besides a binary final answer correctness. Although they recommended to remove 10.3% of the original test instances based on a clinician review, they otherwise treated the rest of the original MedCalc-Bench reference labels as valid. Our work is complementary yet distinct in our focus on developing a phased, physician-in-the-loop stewardship protocol that uses automated triage to amplify the impact of precious physician feedback, ultimately flagging 27.3% of original test instances as likely label errors and independently recomputing their labels. The test instances we updated is more than 2.6 times as theirs. Furthermore, we show the downstream impact of the label noise by showing that training a language model (using RL) on the revised versus original MedCalc-Bench reference labels would lead to a sizable distortion of perceived test accuracy, underscoring the importance of reference label choice.
Since November 2025, the MedCalc-Bench authors have been posting dataset updates, starting with v1.1 and v1.2, [^2] Permanent URLs to (v1.1) and (v1.2) of the MedCalc-Bench dataset, which were posted by a MedCalc-Bench author during November 2025 as two separate data repositories on Hugging Face. on both GitHub and Hugging Face; each version release was presented as a replacement of the original v1.0 dataset published in NeurIPS conference. Then in December 2025, they started a new GitHub repository named “MedCalc-Bench-Verified”, which describes itself as a replacement for v1.2, [^3] https://github.com/nikhilk7153/MedCalc-Bench-Verified. The new repository describes itself as a successor of the original (v1.0) repository at https://github.com/ncbi-nlp/MedCalc-Bench. though the exact sense in which this new release was “verified” relative to the peer-reviewed NeurIPS v1.0 release remains unclear. The latter repository’s content was last revised on January 19, 2026, appearing to be unfinished as of our writing.
@@
Appendix C Section 3.3 Mathematical Derivations
@@
On MedCalc’s full test set ($n=1047$), Phase 2 (§3.2) produces a high-confidence $\hat{y}^{\mathrm{new}}$ label for $887$ instances, in which $286$ are flagged as “Likely Error” ($220$ with $\textsf{rel.err}(\hat{y}^{\mathrm{original}},\hat{y}^{\mathrm{new}})>0.05$ plus $66$ $\mathsf{N/A}$); the remaining 761=1047286761=1047-286 are unflagged. Phase 3 (§3.3) then samples $50$ instances from the $286$ flagged instances for physician recomputation, reporting $\mathrm{sMAPE}(\hat{y}^{\mathrm{original}},y^{\star})=72.7\%$ (95% CI [47.9, 99.1]) vs $\mathrm{sMAPE}(\hat{y}^{\mathrm{new}},y^{\star})=72.7\%$ (95% CI [8.0, 35.9]) against physician gold-standard labels ($y^{\star}$).
@@
Since the unflagged stratum ($n_{\text{unflag}}=761$) is treated as having fixed constant error values ($K$) under our counterfactual assumptions, the variance in the test-wide estimates $\bar{s}$ arises solely from the sampling uncertainty in the flagged stratum ($n_{\text{flag}}=286$). We therefore derive the test-wide confidence intervals by linearly scaling the bootstrap confidence intervals from Table 1.
@@
Appendix D Computing Medical Scores: a Formal Lens
In this paper, “task” and “workflow” are two related but different concepts. A task, we say, is a directive to obtain a verifiable outcome when given an input, without specifying the ideal problem-solving procedure. Given a task, a workflow is the operational process which an agent can execute to complete the task. In the following, we mathematically abstract how medical computing workflows are often operationalized in practice, and explain the state-of-the-art public benchmark environment for evaluating an LLM agent’s capability on medical computing tasks.
Table 2: Mapping Formalism to a Concrete Example: The LACE Score[42]
@@ D.1 Mathematical formalism
    categorical values, ranging from True/False to symptom groups—we call this a set $\mathcal{C}$;
@@
The above formalism abstracts the sequential modularity of a typical medical computing workflow and the graphical structure of information flows: the first stage ($f_{\theta}$) interprets and extracts structured intermediate variables from unstructured data, while the second stage ($g_{\phi}$) performs a rule-based routine to aggregate the intermediate bits into a clinically verifiable and actionable scalar output. Of course, not all end-to-end workflows follow this specific pipeline structure; a system designer may as well eliminate $g_{\phi}$ and use a single call of $f_{\theta}$ to directly output the scalar output. That said, [16] employed the former, structured data generation approach, so understanding this allows us to attribute any deviation in the final output $y$ to a parameter extraction error in $f_{\theta}$, or a formulaic mistake in $g_{\phi}$, or both.
@@ D.2 Current task evaluation standard
To our best knowledge, the only large-scale public benchmark dataset for evaluating an agentic system’s end-to-end score computing capability is MedCalc-Bench, created and released by researchers at NIH/NLM and five US medical schools.[16] Their benchmark task asks an LLM model to read a patient note and answer a medical score computation question, and the model’s final answer $y$ is graded correct if it is within a $\pm 5\%$ margin of a gold label value $\hat{y}^{\mathrm{original}}$ that the benchmark authors synthesized; and in cases where $y$ should be a datetime, exact string match is required. A model’s performance is quantified as the average accuracy on the benchmark’s test data set, which are 1047 instances manually reviewed by those authors.
@@
[16] GPT-4 one-shot 50.9
[28] Llama3-70B 4-bit quantized 40.0
[43] GPT-4o RAG + code 59.6*
[23] GPT-4o multi-agent + tool use 67.7
[47] Qwen2.5-7B RLVR 57.6
[21] DeepSeek-R1 zero-shot 48.1
[21] Llama3-3B SFT + RLVR 52.0
@@
Since June 2024, multiple academic studies have used MedCalc-Bench as an evaluation benchmark or finetuned open-weight models on its train set. Table 3 shows a partial anthology of some test accuracies recently reported. Several survey papers on LLM and healthcare, including [45, 48, 46], cited this benchmark. In addition, MedCalc-Bench is also included in the test suite of MedHELM,[6] a comprehensive framework for evaluating medical LLMs. These references indicate that MedCalc-Bench is an influential benchmark in the field. As all these studies were posted online before October 2025, i.e. before MedCalc-Bench authors had publicly recommended any changes to their dataset version in November 2025, it is fairly certain that their evaluations were based on the original NeurIPS version, i.e. (v1.0).
@@ D.4 Model alignment as policy optimization
We adopt a standard sequential decision-making formalism for model alignment, following the RL literature on policy-gradient methods and recent work on LLM post-training. [^5] See, e.g. [44, 38, 33, 17, 29, 2, 13, 19]. In this view, the interaction between an LLM agent and its digital environment is a Markov decision process (MDP) $(\mathcal{S},\mathcal{A},R,P,\mu)$: at each discrete time $t=1,\dots,T$ the agent observes a state $s_{t}\in\mathcal{S}$, samples an action $a_{t}\in\mathcal{A}$ from a stochastic policy $\pi_{\theta}(a_{t}\mid s_{t})$, receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$, and the environment transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. A trajectory (i.e. a token sequence for an LLM) is
@@ Appendix E Controlled RL Experiment Details
Our goal is to isolate the causal effect of label choice on an LLM’s model alignment when RL is used. To this end, we run two GRPO training experiments on the same base model that are identical in all respects except for the labels used to define the training reward. In the original-label arm, training rewards (as assigned by (5)) are computed against MedCalc-Bench’s original labels $\hat{y}^{\mathrm{original}}$; in the new-label arm, training rewards are computed against our corrected labels $\hat{y}^{\mathrm{new}}$. At test time, the 0-1 accuracy of both models are always graded against $\hat{y}^{\mathrm{new}}$, which our Phase 2 study (§3.3) found to be more clinically accurate than $\hat{y}^{\mathrm{original}}$.
@@
Both RL runs are trained on the same subset of 4,5934{,}593 MedCalc-Bench train instances for which Pipeline 2 produced high-confidence $\hat{y}^{\mathrm{new}}$ labels. The held-out test set consists of $887$ instances whose $\hat{y}^{\mathrm{new}}$ labels were also obtained from Pipeline 2 and partially audited by physicians. In both arms, the policy LLM sees exactly the same collection of prompts and calculator types; the only difference is whether the training reward’s $y^{*}$ is instantiated using $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$. The underlying system and user prompts, including chain-of-thought and answer-format instructions, are held fixed across runs; we provide the full prompt template in the Appendix.
@@
In both arms, the trainable policy $\pi_{\theta}$ and the frozen reference policy $\pi_{\text{ref}}$ are initialized from the same Qwen3-8B base model,[31] chosen as a widely used open model for academic post-training. $\pi_{\text{ref}}$, being used for KL regularization, remains fixed throughout training, while $\pi_{\theta}$ is updated iteratively by gradient descent to minimize the GRPO loss in Eq. 9. We use the same updated system prompt (which allows for abstention) and require chain-of-thought style completions with a final scalar answer token; no external tools are used during RL training, and the interaction is purely text-based.
@@
We instantiate the policy objective from §D.4 using the trajectory-level reward $r(\tau)$ in Eq. 5 with $\lambda_{f}=0.1$. For the comparison experiment in §2.6, both runs share the same optimizer and hyperparameters: Adam with learning rate $\text{lr}=10^{-5}$, KL regularization coefficient $\beta=10^{-3}$, and no reward standardization or length normalization to ensure advantage estimation is unbiased.[24] Each GRPO minibatch consists of $256$ distinct $(C,q)$ pairs from the MedCalc training set. We maintain the same random seed so that the minibatch partitions and ordering are identical between the two runs. We also conducted a hyperparameter sensitivity analysis over the learning rates $\{10^{-4},10^{-5},10^{-6}\}$ and KL coefficient $\{10^{-2},10^{-3},10^{-4}\}$.
@@
During training, all trajectories are generated from $\pi_{\theta}$ with a temperature of $1.0$, top-$p=1.0$, and a maximum output length of $1600$ tokens. Moreover, the GRPO group size $G=8$, i.e. $8$ trajectories are sampled i.i.d. from $\pi_{\theta}$ for each $(C,q)$ pair in a minibatch. After training, we evaluate the two policies on the $887$-instance held-out test set, with ths same inference configuraiton as training rollouts. The evaluation metric is the 0-1 test accuracy defined in §3.3: we parse the model’s final scalar score (or abstention token) and compare it against the physician-preferred $\hat{y}^{\mathrm{new}}$ labels, treating abstentions as correct only when $\hat{y}^{\mathrm{new}}$ is NA. All decision rules—prompts, reward shape, decoding, and accuracy metric—are held fixed across the two training runs; only the training labels used inside $r(\tau)$ differ.

### v3

@@ Appendix A LLM Prompts
Phase 2 System Prompt: You will read a user-provided JSON string that contains an anonymized patient note (‘‘Patient Note’’) and a clinical computation question (‘‘Question’’). Your job is to calculate the answer to the question based on the patient note. You have access to a Google Search tool and a code execution tool to perform calculations, if needed. Your operational protocol is as follows: 1. Adopt Expert Persona: Assume the role of a diligent clinician. Your primary responsibility is accuracy and adherence to standard and up-to-date medical guidelines. 2. Identify Appropriate Scoring Rule or Formula: You will be given a patient note and a question that specifies a clinical calculation task (e.g., CHA2DS2-VASc score, MELD score, CURB-65, etc.). You should use Google Search to confirm the current, generally accepted formula, criteria, and/or scoring system for the specified clinical computation question. Prioritize reputable medical sources, such as MDCalc.com and guidelines published by health agencies or hospitals. 3. Systematic Reasoning and Calculation: $\bullet$ Carefully read the patient note and systematically extract every piece of information relevant to the calculation. $\bullet$ Explicitly map the extracted patient data to the components of the scoring system. For example, for a CHA2DS2-VASc score, list each component (Congestive heart failure, Hypertension, Age, etc.) and assign points based on the patient note. $\bullet$ If there are complex arithmetic operations that are not easily performed by hand, you can use the code execution tool to perform the calculations. $\bullet$ Show your step-by-step calculation. Sum the points or apply the formula as required. 4. Handling Missing Information: $\bullet$ If a piece of information required for the calculation is clearly absent from the note and if it cannot be reasonably inferred from the note, explicitly state what is missing in your reasoning, and give your final answer as ‘‘unknown’’ (without quotes). $\bullet$ Do not invent data or make assumptions that are not reasonably supported by the patient note. 5. Final Answer Formatting (Strict): $\bullet$ After your detailed reasoning, you MUST enclose your final answer in <answer> </answer> tags. $\bullet$ The content inside the tags must ONLY be the final numeric value (e.g., 4, 12.5), date, or time. For numeric values, you do not need to round the answer. $\bullet$ As mentioned in #4, if and only if the calculation is impossible due to missing data, the content inside the answer tags must be exactly ‘‘unknown’’ (without quotes).  $\bullet$ DO NOT include any units, variable names, explanations, or any other text inside the <answer> </answer> tags. Phase 2 User Prompt Example: { "Patient Note": "An 87-year-old man was admitted to our hospital for anorexia for several days, high-grade fever from the previous day, and liver dysfunction. Of note, he had a history of hypertension, diabetes mellitus (DM), and angina. Physical examination findings included: clear consciousness; height, 163 cm; weight, 48 kg; blood pressure, 66/40 mmHg; heart rate, 75/min; respiratory rate, 22/min...[Omitted content was present in the actual API call]...causes of death were thus determined as (1) DIC and (2) ALF associated with diffuse large B-cell lymphoma.", "Question": "What is the patient’s Creatinine Clearance using the Cockroft-Gault Equation in terms of mL/min? You should use the patient’s adjusted body weight in kg instead of the patient’s actual body weight if the patient is overweight or obese based on their BMI. If the patient’s BMI is normal, set their adjusted body weight to the minimum of the ideal body and actual weight. If the patient is underweight, please set their adjusted body weight to their actual body weight.", }
@@ Appendix B Concurrent Work
Wang et al. (Wang et al., 2025) reexamined MedCalc-Bench by introducing an LLM-judged, stepwise evaluation pipeline to grade an LLM’s intermediate calculation steps at test time, which can be thought of as providing process-level rewards besides a binary final answer correctness. Although they recommended to remove 10.3% of the original test instances based on a clinician review, they otherwise treated the rest of the original MedCalc-Bench reference labels as valid. Our work is complementary yet distinct in our focus on developing a phased, physician-in-the-loop stewardship protocol that uses automated triage to amplify the impact of precious physician feedback, ultimately flagging 27.3% of original test instances as likely label errors and independently recomputing their labels. The test instances we updated is more than 2.6 times as theirs. Furthermore, we show the downstream impact of the label noise by showing that training a language model (using RL) on the revised versus original MedCalc-Bench reference labels would lead to a sizable distortion of perceived test accuracy, underscoring the importance of reference label choice.
Since November 2025, the MedCalc-Bench authors have been posting dataset updates, starting with v1.1 and v1.2, [^2] Permanent URLs to (v1.1) and (v1.2) of the MedCalc-Bench dataset, which were posted by a MedCalc-Bench author during November 2025 as two separate data repositories on Hugging Face. on both GitHub and Hugging Face; each version release was presented as a replacement of the original v1.0 dataset published in NeurIPS conference. Then in December 2025, they started a new GitHub repository named “MedCalc-Bench-Verified”, which describes itself as a replacement for v1.2, [^3] https://github.com/nikhilk7153/MedCalc-Bench-Verified. The new repository describes itself as a successor of the original (v1.0) repository at https://github.com/ncbi-nlp/MedCalc-Bench. though the exact mechanism in which this new release was “verified” relative to the peer-reviewed NeurIPS v1.0 release remains unspecified. The latter repository’s content was last revised on January 19, 2026, appearing to be unfinished as of our writing.
@@
Appendix C Test-wide Error Estimation and Sources of Disagreement
In Phase 3 (§3.3), we reported sMAPE values comparing the original ($\hat{y}^{\mathrm{original}}$) and recomputed ($\hat{y}^{\mathrm{new}}$) labels against physician ground truth on a 50-instance subsample of the test set. This appendix asks: how does this comparison translate to the entire 1,047-instance test set? Direct measurement is infeasible because producing physician ground truth for the remaining instances would require an order of magnitude more clinician time than was available. We instead derive a conservative estimate from two ingredients.
First, our 50 physician-adjudicated instances were drawn from the 286 cases Phase 2 flagged as likely errors. We assume that the sMAPE values measured on this sample are representative of the full flagged stratum: i.e., that physician disagreement would scale similarly if more flagged cases were adjudicated. This is the only true assumption we make.
Second, on the 761 unflagged test instances, $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ are within $\pm 5\%$ of each other by construction. This is not an assumption but a fact: it follows directly from how Phase 2’s triage rule defines “unflagged.” The implication is that on the unflagged stratum, the two estimators cannot meaningfully disagree on accuracy: if the original label is close to ground truth, the recomputed one must be too, and vice versa. We use this fact to bound the worst-case sMAPE on the unflagged stratum.
Combining these two ingredients yields a stratified test-wide estimate that conservatively favors the original benchmark. The remainder of this appendix formalizes this argument.
@@ Appendix C Test-wide Error Estimation and Sources of Disagreement
On MedCalc’s full test set ($n=1047$), Phase 2 (§3.2) produces a high-confidence $\hat{y}^{\mathrm{new}}$ label for $887$ instances, in which $286$ are flagged as “Likely Error” ($220$ with $\textsf{rel.err}(\hat{y}^{\mathrm{original}},\hat{y}^{\mathrm{new}})>0.05$ plus $66$ $\mathsf{N/A}$); the remaining 761=1047286761=1047-286 are unflagged. Phase 3 (§3.3) then samples $50$ instances from the $286$ flagged instances for physician recomputation, reporting $\mathrm{sMAPE}(\hat{y}^{\mathrm{original}},y^{\star})=72.7\%$ (95% CI [47.9, 99.1]) vs $\mathrm{sMAPE}(\hat{y}^{\mathrm{new}},y^{\star})=20.1\%$ (95% CI [8.0, 35.9]) against physician gold-standard labels ($y^{\star}$).
@@
Since the unflagged stratum ($n_{\text{unflag}}=761$) is treated as having fixed constant error values ($K$) under our counterfactual assumptions, the variance in the test-wide estimates $\bar{s}$ arises solely from the sampling uncertainty in the flagged stratum ($n_{\text{flag}}=286$). We therefore derive the test-wide confidence intervals by linearly scaling the bootstrap confidence intervals from Figure 3.
@@
C.1 Qualitative Sources of Disagreement
Beyond the quantitative argument above, the physicians’ free-text comments collected during Phase 3 surface recurring qualitative reasons why a single “ground truth” label may not exist for many instances. We summarize these patterns below. Retrospective comparison against single-blind physician recomputations suggests that a substantial share of label disagreement is driven by context underspecification: many original $(C,q)$ instances do not contain enough contextual qualifiers to define a unique clinical “ground truth” score. Two patterns recur. First, the benchmark rarely specifies a scoring timepoint. Many notes describe multi-visit trajectories over weeks or months, but the question does not clarify whether the score should be computed at presentation, peak severity, discharge, or another reference time (e.g. rows 138, 334, 341, 509, 569, 739, 761, 785). Second, scoring rules can vary across time and jurisdiction, so the “official” aggregation rule is not uniquely describable by a static benchmark label. For example, Morphine Milligram Equivalent (MME) conversions for methadone depend on which guideline is assumed: pre-2023 CDC guidance used a tiered rule, i.e. a multiplier of 12 for methadone doses >60>60 mg/day (CDC, 2016), whereas the current CDC guideline adopts a flatter multiplier (4.7) while cautioning that conversion becomes unreliable at higher doses (Dowell et al., 2022); still, today’s evidence suggest that some state-level guidelines reflect the older CDC schedule (Maryland Department of Health, 2025; Utah Department of Health and Human Services, 2025). Under such ambiguity, multiple defensible answers exist and a single numeric gold label does not cleanly measure an LLM’s generalizable clinical capability.
Appendix D Technical Background: From Score Computation to Model Training
This appendix collects the formal definitions and derivations that underpin the main text. It is organized in four parts. §D.1 formalizes medical score computation as a two-stage pipeline (feature extraction followed by rule-based aggregation), identifying the points at which labeling errors can enter; Table 1 grounds this formalism in a concrete LACE score example. §D.2 describes MedCalc-Bench’s evaluation protocol and summarizes prior reported results. §D.3 discusses what it means to treat a benchmark label as a point estimate of a latent ground truth, and the implications for inter-rater variability. Finally, §D.4 provides a self-contained derivation of the GRPO algorithm used in our RL experiment, adapted to the on-policy, trajectory-level reward setting relevant to medical score computation.
Table 1: Mapping Formalism to a Concrete Example: The LACE Score (Van Walraven et al., 2010)
@@ D.1 Mathematical formalism
    categorical values, ranging from True/False to symptom groups; we call this a set $\mathcal{C}$;
@@
The above formalism abstracts the sequential modularity of a typical medical computing workflow and the graphical structure of information flows: the first stage ($f_{\theta}$) interprets and extracts structured intermediate variables from unstructured data, while the second stage ($g_{\phi}$) performs a rule-based routine to aggregate the intermediate bits into a clinically verifiable and actionable scalar output. Of course, not all end-to-end workflows follow this specific pipeline structure; a system designer may as well eliminate $g_{\phi}$ and use a single call of $f_{\theta}$ to directly output the scalar output. That said, (Khandekar et al., 2024) employed the former, structured data generation approach, so understanding this allows us to attribute any deviation in the final output $y$ to a parameter extraction error in $f_{\theta}$, or a formulaic mistake in $g_{\phi}$, or both.
@@ D.2 Current task evaluation standard
To our best knowledge, the only large-scale public benchmark dataset for evaluating an agentic system’s end-to-end score computing capability is MedCalc-Bench, created and released by researchers at NIH/NLM and five US medical schools (Khandekar et al., 2024). Their benchmark task asks an LLM model to read a patient note and answer a medical score computation question, and the model’s final answer $y$ is graded correct if it is within a $\pm 5\%$ margin of a gold label value $\hat{y}^{\mathrm{original}}$ that the benchmark authors synthesized; and in cases where $y$ should be a datetime, exact string match is required. A model’s performance is quantified as the average accuracy on the benchmark’s test data set, which are 1047 instances manually reviewed by those authors.
@@
Khandekar et al. (Khandekar et al., 2024) GPT-4 one-shot 50.9
Oh et al. (Oh et al., 2025) Llama3-70B 4-bit quantized 40.0
Bedi et al. (Bedi et al., 2026) o3-mini zero-shot 34.0
Wang et al. (Wang et al., 2025) GPT-4o RAG + code 59.6*
Liu et al. (Liu et al., 2025a) GPT-4o multi-agent + tool use 67.7
Zhang et al. (Zhang et al., 2025) Qwen2.5-7B RLVR 57.6
Lin et al. (Lin et al., 2025) DeepSeek-R1 zero-shot 48.1
Lin et al. (Lin et al., 2025) Llama3-3B SFT + RLVR 52.0
@@
Since June 2024, multiple academic studies have used MedCalc-Bench as an evaluation benchmark or finetuned open-weight models on its train set. Table 2 shows a partial anthology of some test accuracies recently reported. Several survey papers on LLM and healthcare, including (Xie et al., 2024; Zhou et al., 2025; Xu and Sankar, 2025), cited this benchmark. In addition, MedCalc-Bench is also highlighted as a representative quantitative clinical benchmark in the test suite of MedHELM (Bedi et al., 2026), a comprehensive framework for evaluating medical LLMs, recently published on Nature Medicine. These references indicate that MedCalc-Bench is an influential benchmark in the field. As all these studies were posted online before October 2025, i.e. before MedCalc-Bench authors had publicly recommended any changes to their dataset version in November 2025, it is fairly certain that their evaluations were based on the original NeurIPS version, i.e. (v1.0).
@@ D.4 Model alignment as policy optimization
We adopt a standard sequential decision-making formalism for model alignment, following the RL literature on policy-gradient methods and recent work on LLM post-training. [^5] See, e.g. (Williams, 1992; Sutton et al., 1999; Schulman et al., 2016; Kool et al., 2019; Ouyang et al., 2022; Ahmadian et al., 2024; Guo et al., 2025; Kumar et al., 2025). In this view, the interaction between an LLM agent and its digital environment is a Markov decision process (MDP) $(\mathcal{S},\mathcal{A},R,P,\mu)$: at each discrete time $t=1,\dots,T$ the agent observes a state $s_{t}\in\mathcal{S}$, samples an action $a_{t}\in\mathcal{A}$ from a stochastic policy $\pi_{\theta}(a_{t}\mid s_{t})$, receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$, and the environment transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. A trajectory (i.e. a token sequence for an LLM) is
@@ Appendix E Controlled RL Experiment Details
Our goal is to isolate the effect of label choice on an LLM’s alignment when RL is used. We run two GRPO training experiments on the same base model that are identical in all respects except for the labels used to define the training reward. In the original-label arm, training rewards (as assigned by (5)) are computed against MedCalc-Bench’s original labels $\hat{y}^{\mathrm{original}}$; in the new-label arm, training rewards are computed against our recomputed labels $\hat{y}^{\mathrm{new}}$.
Evaluation strategy.
We evaluate the two trained models along three dimensions, each reported in a different section:
  1. 1.
    Primary evaluation against physician ground truth ($y^{\star}$, $n=50$). Both models are evaluated on the 50 physician-labeled test instances from Phase 3, where we have direct access to clinician-computed labels. This is our primary measure of whether label quality during training translates into clinically meaningful differences. The main result is reported in the Results section (§3.4).
  2. 2.
    Full training dynamics on the 887-instance test set (graded against $\hat{y}^{\mathrm{new}}$). We track accuracy over the course of training on the larger held-out set, using $\hat{y}^{\mathrm{new}}$ as a proxy for ground truth. This provides a higher-powered view of the performance gap. Details are in Appendix §E.1; a summary appears in §3.4.
  3. 3.
    Generalization to additional medical evaluation sets. We test both checkpoints on MedQA and two recent revisions of MedCalc-Bench to assess whether the training-label effect transfers to related medical tasks. Details are in Appendix §E.2; a summary appears in §3.4.
@@
Both RL runs are trained on the same subset of 4,5934{,}593 MedCalc-Bench train instances for which Phase 2 produced high-confidence $\hat{y}^{\mathrm{new}}$ labels. The held-out test set consists of $887$ instances whose $\hat{y}^{\mathrm{new}}$ labels were also obtained from Phase 2 and partially audited by physicians. In both arms, the policy LLM sees exactly the same collection of prompts and calculator types; the only difference is whether the training reward’s $y^{*}$ is instantiated using $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$. The underlying system and user prompts, including chain-of-thought and answer-format instructions, are held fixed across runs; we provide the full prompt template in Appendix §A.
@@
In both arms, the trainable policy $\pi_{\theta}$ and the frozen reference policy $\pi_{\text{ref}}$ are initialized from the same Qwen3-8B base model (Qwen, 2025), chosen as a widely used open model for academic post-training. $\pi_{\text{ref}}$, being used for KL regularization, remains fixed throughout training, while $\pi_{\theta}$ is updated iteratively by gradient descent to minimize the GRPO loss in Eq. 9. We use the same updated system prompt (which allows for abstention) and require chain-of-thought style completions with a final scalar answer token; no external tools are used during RL training, and the interaction is purely text-based.
@@
We instantiate the policy objective from §D.4 using the trajectory-level reward $r(\tau)$ in Eq. 5 with $\lambda_{f}=0.1$. For the comparison experiment in §2.6, both runs share the same optimizer and hyperparameters: Adam with learning rate $\text{lr}=10^{-5}$, KL regularization coefficient $\beta=10^{-3}$, and no reward standardization or length normalization to ensure advantage estimation is unbiased (Liu et al., 2025b). Each GRPO minibatch consists of $256$ distinct $(C,q)$ pairs from the MedCalc training set. We maintain the same random seed so that the minibatch partitions and ordering are identical between the two runs. We also conducted a hyperparameter sensitivity analysis over the learning rates $\{10^{-4},10^{-5},10^{-6}\}$ and KL coefficient $\{10^{-2},10^{-3},10^{-4}\}$.
@@
During training, all trajectories are generated from $\pi_{\theta}$ with a temperature of $1.0$, top-$p=1.0$, and a maximum output length of $1600$ tokens. Moreover, the GRPO group size $G=8$, i.e. $8$ trajectories are sampled i.i.d. from $\pi_{\theta}$ for each $(C,q)$ pair in a minibatch. After training, we evaluate the two policies on the $887$-instance held-out test set, with ths same inference configuraiton as training rollouts. The evaluation metric is the 0-1 test accuracy defined in §3.3: we parse the model’s final scalar score (or abstention token) and compare it against the physician-preferred $\hat{y}^{\mathrm{new}}$ labels, treating abstentions as correct only when $\hat{y}^{\mathrm{new}}$ is NA. All decision rules (prompts, reward shape, decoding, and accuracy metric) are held fixed across the two training runs; only the training labels used inside $r(\tau)$ differ.